I need this immediately for my H.W : The total surfac area of a cone is 616 sq cm.Its slant height i.e. length is equal to three times the radius of the base.Find its radius and slant height.
Please please please send the answer fast...

Answer by venugopalramana(3286)

I need this immediately for my H.W : The total surfac area of a cone is 616 sq cm.Its slant height i.e. length is equal to three times the radius of the base.Find its radius and slant height.
Please please please send the answer fast...
TSA=PI*R^2+PI*R*L=616
LET R BE THE RADIUS.SO SLANT HT.=L=3R
PI(R^2+R*3R)=616
PI*4R^2=616
R^2=616/(4PI)=616*7=/(4*22)=7*7
R=7
SLANT HT.=3*7=21